OpenStudy (priyar):

Two soap bubbles coalese to form a single bubble . If V is the subsequent change in volume of contained air and S the change in total surface area , T is the surface tension and P is the atmospheric pressure . Find a relation between P,V,S,T

OpenStudy (priyar):

@ParthKohli @ganeshie8

OpenStudy (priyar):

but how to bring in Patm ?

OpenStudy (priyar):

@whpalmer4 @imqwerty @rvc

OpenStudy (priyar):

@chmvijay can u help?

OpenStudy (priyar):

@Awolflover1

OpenStudy (shamim):

If 2 bubbles make a single bubble, then the total surface area will b decrease. Right?

OpenStudy (shamim):

And energy will b released

OpenStudy (shamim):

If u break a single bubble into many bubbles , then the surface area will b increase nd energy needed to do so

OpenStudy (shamim):

In this case Energy release=TS

OpenStudy (shamim):

Here S=increase of surface area

OpenStudy (shamim):

T=surface tension

OpenStudy (priyar):

yes u r right...

OpenStudy (priyar):

i have calculated S=2r1r2

OpenStudy (priyar):

* S=-2r1r2 right?

OpenStudy (shamim):

I dont know how S=2r1r2

OpenStudy (priyar):

wait i will check again..now i think its wrong..

OpenStudy (priyar):

r1^3+ r2^3 =R^3

OpenStudy (priyar):

but if i go by this method and find V and S its very time consuming and lengthy...is there any short method?

OpenStudy (priyar):

hey wont V be zero?

OpenStudy (priyar):

the volume of two bubbles will be equal to the volume of the bigger bubble right?

OpenStudy (shamim):

Yes u r right! Volume of 2 bubbles = volume of bigger bubble

OpenStudy (shamim):

Nd volume of a sphere=(4/3)πr^3

OpenStudy (priyar):

so how can we find the relation? Subsequent change in volume is zero..??

OpenStudy (shamim):

So we can write (4/3)πr1^3+(4/3)πr2^3=(4/3)πR^3 Right?

OpenStudy (priyar):

this what i used to find the relation..(i have written it above..did u see?)

OpenStudy (shamim):

U wrote r1^3+r2^3=R^3 Which wrong

OpenStudy (priyar):

why?

OpenStudy (priyar):

why is it wrong?

OpenStudy (shamim):

Sorry

OpenStudy (shamim):

U r also correct !!

OpenStudy (priyar):

its ok!

OpenStudy (shamim):

Anyway i dont know that relation berween those

OpenStudy (priyar):

but u didn't my previous doubt.....

OpenStudy (priyar):

V=0 right?

OpenStudy (priyar):

@shamim ?

OpenStudy (shamim):

Yes change of volume = 0

OpenStudy (priyar):

then how can we find a relation bet V,T,S,P?? if one is zero?

OpenStudy (priyar):

@ganeshie8 @samigupta8 @tkhunny @agent0smith

OpenStudy (samigupta8):

Okk @priyar i found a way...

OpenStudy (priyar):

great what is it?

OpenStudy (samigupta8):

Luk it's not the volume addition in any way

OpenStudy (samigupta8):

What u have to do is conserve the moles

OpenStudy (priyar):

isn't the volume same?

OpenStudy (samigupta8):

n1+n2=N N1 ,N2 r the moles of air in 1 n 2nd soap bubble N is that of bigger bubble

OpenStudy (priyar):

oh ok..

OpenStudy (samigupta8):

U got whati want to say priyar

OpenStudy (priyar):

ya..u r conserving the moles right? but i have a doubt..why won't the volume be same?

OpenStudy (samigupta8):

U myt have done gaseous state ....what do u do over there do u add up the volumes..of gas no naa...den here also d same thing rules

OpenStudy (samigupta8):

Is d ans 3PV+ 4ST=0

OpenStudy (priyar):

but in fluids we do that always..

OpenStudy (samigupta8):

Where do we use dis thing in fluids

OpenStudy (priyar):

usually to find the radius of the bigger bubble formed by two smaller bubbles we use the fact that the volume remains the same..right?

OpenStudy (priyar):

btw the answer is correct! how did u get it?

OpenStudy (samigupta8):

The same way i told u ...u need to work upon it...there will be like P1V1 + P2V2 =PV

OpenStudy (samigupta8):

From conservation of moles....

OpenStudy (samigupta8):

Now put in the value of P1 as Po+4T/R1 n same put in the value of P2 n P also

OpenStudy (samigupta8):

Plug in the value of volumes also as they r spheres only....4/3πr^3

OpenStudy (priyar):

Po+4T/R1 ?

OpenStudy (samigupta8):

Excess pressure

OpenStudy (priyar):

what is ur Po and R1? first tell me the notations

OpenStudy (samigupta8):

Po is atmospheric pressure n R1 is radius of frst bubble

OpenStudy (samigupta8):

R u in 11th grade?

OpenStudy (samigupta8):

@priay did u get it?

OpenStudy (priyar):

sorry.. @samigupta8 was seeing my math doubt.. (P+aT/R1)V1+(P+4T/R2)V2=(P+4T/R)V is this correct?

OpenStudy (priyar):

how to proceed?

OpenStudy (priyar):

(P+4T/R1)V1+(P+4T/R2)V2=(P+4T/R)V how to proceed further how to eliminate V1, V2.. @samigupta8..

OpenStudy (samigupta8):

V1 is 4/3πr1^3

OpenStudy (priyar):

but won't it be lengthy..wait i'll do..

OpenStudy (samigupta8):

Same u cn find out V2

OpenStudy (samigupta8):

It is definitely not lengthy i solved it jst calculation based...

OpenStudy (samigupta8):

N pls dpn't cut 4/3 π on both lhs n rhs sides

OpenStudy (samigupta8):

Don't*

OpenStudy (priyar):

oh...i just did that!

OpenStudy (priyar):

(P+4T/R1)4/3 pi R1^3+(P+4T/R2)4/3 pi R2^3=(P+4T/R)4/3piR^3

OpenStudy (samigupta8):

Okk now open the brackets n solve it

OpenStudy (samigupta8):

PoV1 + 16/3πTR1^2 types trms will form

OpenStudy (priyar):

oh then se substitute 4piR1^2 as S1 right?

OpenStudy (samigupta8):

Correct u made it

OpenStudy (priyar):

yay got it!

OpenStudy (priyar):

Thanks..!

OpenStudy (priyar):

@ParthKohli

OpenStudy (priyar):

here won't V be zero??

OpenStudy (samigupta8):

Hey again u r having d same doubt...okk...luk in d question itself it says that V is volume change then why r u noy considering it....

OpenStudy (priyar):

no..that was for parthkohli.. regarding another Q that was related with equating volume..